Greenhab Report – April 27th

Crew 211 GreenHab Report – 27-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 3

Shade cloth: 80%

Average temperatures:

Low: 14°C
High: 46°C

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons

Water in Blue Tank: 90 gallons

Time(s) of watering for crops: 11:00

Changes to crops: All plants look healthy. Green chard, cucumber (03/20) and romaine lettuce grow very well. Tomatoes (from plants that we brought), carrots and radishes are ready to be harvested. Onion plants are emerging. More flowers are visible on the tomato plant of the GreenHab.

Narrative: Still hot today. No harvest. I’m still watering each morning the bean and tomato seeds of my experiment placed in the GreenHab.

Harvest: None

Support/supplies needed: None

Science Report – April 27th

Science Report 27 April 2019
Crew 211 – UCL to Mars
Submitted by Science Officer Maxime Bernard

1) Commander Carl Dahlqvist

He finished his code yesterday and he still testing it today.

He also tested the hardware side of the main module as well as mounted and tested 4 of the 7 secondary modules.

2) XO Officer

Finishing and testing the trilateration algorithm of his ultra-wide bandwidth system.

3) HSO Officer Benjamin Flasse

Every morning and every evening, the HSO still gather data about the measurements, weight, brain awakeness, tension and pulse of the crew.

He has now enough data to start analyzing his CFFF-cortical arousal graphics.

4) Crew Astronomer Eleonore Lieffrig

Working on the density probability via the Direct Fourier transform method while coding an algorithm that should allow her to verify the non-dependency of the colloidal particles Brownian motion with regards to the gravity.

5) Crew Engineer Julien Amalaberque

He studied the performance and efficiency of simultaneous location and mapping algorithm depending on the resolution of video streams coming from Intel RealSence D435i camera.

6) Crew Biologist Chloé Peduzzi

Spent the day preparing her Petri dish.

7) Crew Journalist Nathan Pechon

We had a new meeting to talk about the value of life on Mars and how to deal with criminality.

He then worked on his own on different law articles that resulted from in what was said the previous talks. He is glad these debates aren’t sterile but on the contrary, very constructive and animated.

8) Crew Scientist Maxime Bernard

Keeping up with the data acquisition. Some issues are already appearing with the counting statistic.

Operations Report – April 27th

Crew 211 Operations Report 27-APR-2019

SOL: 6

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 12:00
On (last night): 6:30 pm
Off (this morning): 6:30 am

Solar— SOC 53 % at 19:11 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~150 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 90 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 15 gallons

Toilet tank emptied: no

Deimos rover used: no (out of order)
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: yes
Hours: 89.8h
Beginning charge: 100
Ending charge: 44
Currently charging: yes

Opportunity rover used: no (out of order)
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: yes
Hours: 91.7h
Beginning charge: 100
Ending charge: 47
Currently charging: yes

i. Notes on rovers: Using an extension cable, rovers can now again be charged directly at the MDRS (not at the RAM but near the Hab).

Reason for use: EVA

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support:

* The levels of diesel and propane are absolutely not moving
* Dr. Rupert brought us the first refill of our food with some flour, rice, cereals.. as well as some medical supplies and we are really thankful.

She indicated that we would receive water soon which is also great news.